Q1:Suppose that a worker in Country A can produce either 25 bananas or 5 tomatoes each year. Country A has 200 workers. Suppose

Q2:Tom and Jerry have two tasks to do all day make dishes and build fences. If Tom spends all day making dishes, he will make 16
Answer #1

1. specialize in tomatoes because their opportunity cost of tomatoes is lower than country A's.
(Opportunity cost of 1 tomato in country A = Total bananas given up/Total tomatoes produced = 25/5 = 5 bananas
Opportunity cost of 1 tomato in country B = Total bananas given up/Total tomatoes produced = 18/6 = 3 bananas
So, Country B has comparative advantage in tomatoes because of lower opportunity cost. Thus, country B will benefit if they specialize in tomatoes.)

2. could both benefit from trade because comparative advantage exists.
(Opportunity cost of 1 fence for Tom = Total dishes given up/Total fence built = 16/4 = 4 dishes
Opportunity cost of 1 fence for Jerry = Total dishes given up/Total fence built = 14/7 = 2 dishes
So, Jerry has comparative advantage in fence because of lower opportunity cost and thus they can both gain from trade because comparative advantage exists.)